George has 50 milliliters of a 10 percent acid solution. How many milliliters of 45 percent acid solution should he add to create a 35 percent acid solution?
Answer by checkley77(12844)

.45x+.10*50=.35(50+x)
.45x+5=17.5+.35x
.45x-.35x=17.5-5
.10x=12.5
x=12.5/.10
x=125 ml.of 45% solution is used.
Proof:
.45*125+.10*50=.35(50+125)
56.25+5=.35*175
61.25=61.25